PS-IS-115/RC

 

MINUTES OF THE PS-ISOLDE COORDINATION MEETING

held on 29 July 2002

 

 

Present:

R. Catherall, B. Crepieux, D. Carminati, A. Dorsival, Sven de Man, V. Fedosseev, S. Meunier, J. Parra-Lopez

 

ISOLDE at URL: http://isolde.cern.ch

Consolidation projects at URL:

http://isolde-consolidation.web.cern.ch/isolde-consolidation/

PSMB meetings at URL:

http://cern.web.cern.ch/CERN/Divisions/PS/Meetings/PSMB.html

 

1. HRS:

       The Pb/Fr run with the ThC target was stopped at 08:30. A few measurements were taken in the evening during the radioactive cooling period. UC target #204 to be placed on the separator on Thursday morning (01-08-02).

 

2. GPS:

       During the weekend it was reported that the quadrupole GPS.QP40 (front end) was working incorrectly. Further investigation revealed that the cable from the power supply to the quadrupole was damaged. This was replaced by an existing spare cable.

      

3. LIS

The water problem caused a breakdown this weekend. Valentine suggests that there should be a laser technical stop every 100 hours in order to allow cleaning of the water pipes. He also inquires if a more suitable solution could be found before the shutdown period.

 

Schedule changes

       The Mg run on the GPS has been cancelled.

       The Pb/Fr run was prolonged by one day thus pushing back the target change to Thursday morning. The GPS Mn run should stop for no more than 2 hours to allow this target change.

 

Beam lines

       The cable for CA0.DE60 broke. It has now been repaired (connector changed) by Julien.

       There is a fault with the pumping sequence of section RC0. The vacuum group intends to carry out a leak detection test on this section.

       The storage of equipment around the experiments and amongst the beam lines makes access somewhat difficult. It would be most appreciated if all persons working in the hall could make an effort to store their material in an appropriate manner and remove any redundant material.

 

Vacuum.

       S. Meunier reports that Leybold, under guarantee, will supply a new turbo pump in 2 weeks to replace the broken turbo on the GPS frontend. They would eventually (under proper conditions) like to examine the broken pump.

A spare NT20 controller has been ordered to allow quick testing of the front end pumping system whilst retaining compatibility with the vacuum simatics.

       Various o rings have been ordered for the faraday cups. C. Bal has constructed a vacuum chamber to facilitate the testing of faraday cups and scanners.

      

Chemical laboratory and workshops.

       All users of the above are kindly reminded to leave them as clean and tidy as they would wish to find them. All chemical products removed from the chemical store should be noted in the newly implemented CS logbook and should be returned after use.

 

Other

       Sven de Man is the PS linkman for the integration of EDMS in the PS.

 

Absences

 G. Focker from 20.07 to 18.08

 U. Georg from 26.07 to 31.07

 T. Giles from 29.07 to 02.07

 M. Lindroos from 16.07 to 05.08 (replaced by R. Catherall from 22.07 to

05.08)

 J. Lettry from 22.07 to 10.08 (replaced by R. Catherall)

 T. Nilsson from 11.07 to 31.07 (replaced by F. Herfurth from

 20.07 to 28.07, replaced by Marc Dietrich from 29.07 to 31.07)

 

Next meeting:

Tuesday 6 August 2002 at 11.00 in the EP Conference Room A, building

304-1-001.